OpenStudy (anonymous):

A plane is defined by the equation: x=4y+2z=16. Find two vectors parallel to the plane.

OpenStudy (anonymous):

do you know the rule of parallelism?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

Two vectors and are parallel if their cross product is zero, i.e., .u*v=0

OpenStudy (anonymous):

Okay do what do i do :S @SerikMB

OpenStudy (anonymous):

do you have exact answers?

OpenStudy (anonymous):

one of them may probable be y=-1/2x+2z

OpenStudy (anonymous):

Like what do i do to come up with that answer ?

OpenStudy (anonymous):

when two equations are parallel they have the same slopes

OpenStudy (anonymous):

as it says parallel the equation is linear so it's form is y=kx+b , but here instead of k it will be the slope

OpenStudy (anonymous):

then as vectors usually start at (0,0) point

OpenStudy (anonymous):

you can substitue it into the formula \[Y_{1}-Y _{2}=m(X _{1}-X _{2})\]

OpenStudy (anonymous):

where m is the slope

OpenStudy (anonymous):

given this equation: x=4y+2z=16.

OpenStudy (anonymous):

yea it's just 3 dimensional

OpenStudy (anonymous):

im still not understanding this